Backaround An adult male died suddenly. Postmortem investigation did not identify any gross or historical lesions. Forensic investigation identified strychnine in several tissues. The presence of strychnine at any concentration is significant since it is not a normal body constituent. Case Facts Male -7okg Lethal dose img/kg, BW Individual found dead, rigor mortis had just developed - Tissue concentrations o Blood 2.3mg/L o Liver 14mg/kg o Urine 12.7mg/l Sausage in the stomach contents contained high concentrations of strychnine The person was found dead (recent death) at noon Questions 1. Calculate the total dose of strychnine consumed (mg). C4) 2. Assume the sausage contained the strychnine and the person ingested o.3kg of sausage. What was the amount of strychnine in the sausage (mg)? Based on the answer in # calculation needed. 2, what type/formulation of strychnine was used? A (3) 3, 4. Estimate the time of exposure to strychnine prior to noon. 5. If the urine concentration was substantially lower for example < amg/L, but other L3) tissues remained high, what does this fact indicate in terms of dose, source, organ (5) function and time to death?

Can you please answer part 1, 2 and 3. This is a question related to pharmacokinetics and toxicology. Please show the detailed answer. thank you
Answer #1

1. The total dose of strychnine consumed can be calculated from the amount of strychnine found in:-

Blood - 2.3mg

Liver - 14mg /kg ( liver weight approx = 1.5 kg of an adult)

= 14× 1.5 = 21mg

Urine = 12.7 mg

Total = 2.3 + 21+ 12.7

= 36 mg

2. Amount of sausage ingested = 0.3 kg

Amount of styrchnine present in 0.3 kg of sausage =36 / 0.3

= 120 mg